015
The speed of sound in Rubber, butyl is 1830 m/s. If its density is 1.35 g/cm^3 then its elastic modulus would be:
1
Expert's answer
2021-09-28T11:14:21-0400

We know that


"c=\sqrt{\frac E\rho},\\\space\\
E=c^2\rho=1830^2·1.35=4520000\text{ Pa}."